OpenStudy (anonymous):

solutions to (2x+3)^2=10?

OpenStudy (anonymous):

x=(-3+-sqrt(10))/2

OpenStudy (anonymous):

-squareroot 10-3/2 and pos. squrroot

OpenStudy (anonymous):

\[(2x+3)^2=10\] \[2x+3\pm\sqrt{10}\] \[2x=-3\pm \sqrt{10}\] \[x=\frac{-3 \pm\sqrt{10}}{2}\]